OpenStudy (firejay5):

Find the value of each variable. If your answer is NOT an integer, leave it in simplest radical form. Show work to get medal! :D My Textbook has work shown for these types of problems.

OpenStudy (firejay5):



OpenStudy (firejay5):

@e.mccormick

OpenStudy (anonymous):

you can use your trigonometry here SOH CAH TOA \(sin\ \theta=\frac{opposite}{adjacent}\ \cos\ \theta=\frac{adjacent}{hypotenuse}\ \tan \theta=\frac{opposite}{adjacent}\)

OpenStudy (firejay5):

this is a 30, 60, 90 Special Right Triangle @Data_LG2

OpenStudy (anonymous):

okay, then use the formula \(SL=1/2H\\LL=\frac{\sqrt{3}}{2}H\\LL=SL \sqrt{3}\) where SL=short leg, LL=long leg, H=hypotenuse